# Demagnetization in micromagnetics: magnetostatic self-interactions of bulk chiral magnetic skyrmions

### v1

@@ V The Euler-Lagrange equations
$\frac{\textup{d}}{\textup{d}t}\bigg{|}_{t=0}E_{\textup{exch}}(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\bm{\varepsilon}\cdot\Delta\mathbf{n}$ (37)
$\frac{\textup{d}}{\textup{d}t}\bigg{|}_{t=0}E_{\textup{DMI}}(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\bm{\varepsilon}\cdot(-2\mathbf{d}_{i}\times\partial_{i}\mathbf{n})$ (38)
$\frac{\textup{d}}{\textup{d}t}\bigg{|}_{t=0}E_{\textup{pot}}(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\bm{\varepsilon}\cdot(0,0,-h-2Kn_{3}).$ (39)
@@
Denote by $\psi_{t}$ the unique solution of (29) with source $-{\rm div}\,\mathbf{n}_{t}$ decaying to 00 at infinity, and $\dot{\psi}=\partial_{t}\psi_{t}|_{t=0}$. It is important to note that, while $\bm{\varepsilon}$ has compact support, neither $\psi=\psi_{0}$ nor $\dot{\psi}$ do: as argued above, they are $1/r$ localized. Care must be taken with boundary terms when computing the variation of $E_{\textup{DDI}}$, therefore. The variation of $E_{\textup{DDI}}$ induced by $\mathbf{n}_{t}$ is
@@
=\,= $\frac{1}{\mu}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\psi\Delta\dot{\psi}$
@@
=\,= $\frac{1}{\mu}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\psi\Delta\dot{\psi}$
@@
=\,= $\frac{1}{\mu}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\psi\Delta\dot{\psi}$ (40)
@@
=\,= $\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,{\bm{\varepsilon}}\cdot\bm{\nabla}\psi$
@@
=\,= $\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\bm{\varepsilon}\cdot\bm{\nabla}\psi$ (42)
@@
$\frac{\textup{d}}{\textup{d}t}\bigg{|}_{t=0}E(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\bm{\varepsilon}\cdot(\Delta\mathbf{n}-2\mathbf{d}_{i}\times\partial_{i}\mathbf{n}$ (43)
@@ VI Numerical method
We now detail our method for finding magnetization fields $\mathbf{n}$ that simultaneously minimize the energy (31) and solve Poisson’s equation (29). This method is implemented for NVIDIA CUDA architecture and is adapted from a similar non-local method developed to determine (nuclear) skyrmion crystals stabilized by $\omega$-mesons [27, 28]. Similar methods have also recently been developed to study knot solitons in an extended version of the standard model of particle physics [29] and the back-reaction of the Coulomb force in an extension of the Skyrme model [30]. We follow the same methodology, separating our algorithm into two stages, detailed as follows.
@@ VII Results
We choose to use the micromagnetic parameters $\mathcal{D}=4\,\textup{mJm}^{-2}$, $J=40\,\textup{pJm}^{-1}$, $M_{s}=580\,\textup{kAm}^{-1}$ and $K_{m}=0.8\,\textup{MJm}^{-3}$, in the absence of an external magnetic field, $B_{\textup{ext}}=0\,\textup{T}$. These are the sort of parameters you would use to model something like CoFeB [31]. This gives us the dimensionless coupling constants $K=2$, $\mu=1.057$ and $h=0$. It also defines a length scale of $L_{0}=J/\mathcal{D}=10\,\textup{nm}$ and an energy scale of $E_{0}=J^{2}/\mathcal{D}=4\times 10^{-19}\,\textup{J}$. For simulations involving isolated magnetic skyrmions, all numerics carried out were simulated on a dimensionless meshgrid of size $10\times 10$ with $1024^{2}$ lattice sites and vacuum boundary conditions. Physically, this corresponds to a $100\textup{nm}\times 100\textup{nm}$ grid with $0.098\textup{nm}\times 0.098\textup{nm}$ lattice spacing.
@@ VII.2 Magnetic skyrmion crystals
Next, we investigate the DDI effect on magnetically ordered crystals. In absence of the DDI, there is no optimal lattice geometry – the skyrmions are repulsive and want to be infinitely separated. Our method is simpler than that of [32], we do not vary over the lattice geometry, instead we restrict our geometry to that of two lattice types, an equianharmonic (hexagonal) lattice and a lemniscatic (square) lattice.. That is, we construct our unit cells to be of size $L\times\sqrt{3}L$ and $L\times L$, and we vary the lattice parameter $L$. Then we construct a two (anti)skyrmion initial configuration, consisting of two separated (anti)skyrmions. This is carried out using a product ansatz $u=u_{1}+u_{2}$ in the $\mathbb{C}P^{1}$ formalism. In all cases, as $L\rightarrow\infty$, the energy per unit topological charge approaches that of the isolated single magnetic skyrmion.
@@
In absence of the DDI, there does not exist an optimal lattice size $L_{*}$ which minimizes the energy $E(L)$. The energy is minimized for two infinitely separated skyrmions/antiskyrmions. Consideration of the DDI breaks the energy degeneracy of the crystal configurations, resulting in separate $E$ against $L$ curves for the three DMI terms considered. Qualitatively, the results are similar to that of when no dipolar interaction is considered for the Rashba and Dresselhaus DMIs. That is, we find that the associated skyrmions maintain their repulsive behaviour and want to be infinitely separated, but the Rashba crystals have higher energy than the Dresselhaus ones. Another key difference is that the energy of the Rashba lattice asymptotically approaches the energy of two well-separated Néel skyrmions at a smaller lattice parameter $L$ than the Dresselhaus lattice does to two Bloch skyrmions. This can be seen in Figures 3 and 4. This is most likely due to the fact that the DDI slightly shrinks the size of the Néel skyrmion, making it more localized than the Bloch skyrmion. Even at low lattice parameter $L$, the DDI had minimal effect on the crystalline structure of Dresselhaus Bloch skyrmions, making it indistinguishable from the case without the DDI.
@@
However, a rather striking outcome for the Heusler antiskyrmion lattice is observed. First of all, the antiskyrmions prefer a square lattice arrangement over a hexagonal one at all lattice sizes $L$. Furthermore, it introduces an attractive force such that there does exist an optimal lattice size $L_{*}$ that minimizes the energy $E(L)$. As the lattice size is increased, the energy of the Heusler lattice asymptotically approaches the energy of two infinitely separated antiskyrmions from below, meaning it has positive binding energy $E_{\textup{bind}}=2E_{1}-E_{\textup{latt}}>0$. By contrast, the Rashba and Dresselhaus lattices approach their respective asymptotic energies from above, that is, they have negative binding energies $2E_{1}<E_{\textup{latt}}$. So, it is energetically preferable for the Heusler antiskyrmions to form a crystalline structure, whereas it is preferable for Bloch and Néel skyrmions to be infinitely separated. The results of the effect of the dipolar interaction on the magnetic skyrmion crystals are summarized in Figures 3 and 4.
@@
Figure 4: The effect of the dipolar interaction on square magnetic skyrmion crystals as a function of the lattice size $L$. The unit cell used is of size $L\times L$, which favours a lemniscatic lattice. Initially, two magnetic skyrmions are place in a square arrangement and then are relaxed using the numerical algorithm detailed in the text. It can be seen that including the DDI and its back-reaction lifts the energy degeneracy, resulting in distinct $E(L)$ curves. The Rashba (Néel) and Dresselhaus (Bloch) skyrmion lattices maintain a hexagonal arrangement, whereas the Heusler antiskyrmions relax into one of a square form. An example Heusler antiskyrmion lattice is shown at small $L=20\,\textup{nm}$ and a Rashba Néel skyrmion lattice is shown at a higher $L=50\,\textup{nm}$. Within these subplots is the magnetic charge density $\rho(x,y)$ and the magnetization $\mathbf{n}(x,y)$. Also plotted is the asymptotic energy of two well separated skyrmions for each DMI.
VII.3 Lattice criticality
@@
Our numerical scheme produces, for a given fixed period lattice
@@
and topological degree, the minimizer among all fields $\mathbf{n}:\mathbb{R}^{2}\rightarrow S^{2}$ periodic with respect to $\Lambda$
@@
$\left.+\frac{1}{2\mu}\psi\Delta\psi\right\}$ (50)
where $\Sigma=\mathbb{R}^{2}/\Lambda\equiv T^{2}$ is the torus with period lattice $\Lambda$ (or, in more elementary terms, its unit cell, that is the parallelogram spanned by the vectors $\mathbf{v}_{1},\mathbf{v}_{2}$). We have left the potential terms $V(\mathbf{n})$ arbitrary, since our argument does not rely on the specific potential (4) used in this paper. One should note that such a minimizer is likely to exist for any choice of period lattice $\Lambda$. If this solution is to have any physical significance, it should be a local minimizer of $E_{\textup{cell}}$ with respect to variations of the period lattice $\Lambda$ also. The numerical scheme already checks for criticality with respect to isotropic dilations of $\Lambda$, i.e. $\mathbf{v}_{1}(t)=e^{t}\mathbf{v}_{1}$, $\mathbf{v}_{2}(t)=e^{t}\mathbf{v}_{2}$. It assumes, however, that the energetically optimal shape of the lattice is either equianharmonic or square. Previous work on soliton crystals in other systems [33, 20] shows that this assumption is often ill founded. In this section we derive the conditions for criticality with respect to all variations of the period lattice, adapting ideas introduced in [34].
@@
$-(\partial_{1}^{2}+\partial_{2}^{2})\widetilde{\psi}=-\mu(\partial_{1}\widetilde{\mathbf{n}}+\partial_{2}\widetilde{\mathbf{n}}_{2}).$ (51)
@@
and $\widehat{\psi}:=\psi\circ L:\Sigma\rightarrow{\mathbb{R}}$ satisfies Poisson’s equation on $(\Sigma,\widetilde{g})$
@@
$\frac{\textup{d}}{\textup{d}t}\bigg{|}_{t=0}E_{\textup{pot}}^{\textup{cell}}(A(t))$ $=$ $0$ (67)
@@
VIII Conclusion
An applied external magnetic field was also considered, but the results were qualitatively the same for our parameter set. However, an external magnetic field in the absence of anisotropies was not considered in this study. For this case, as the applied magnetic field strength is increased, one would expect to see the transition from the spin-stripe domain to a lattice of skyrmions as found in [14].
This method could be adapted to the analogue problem of skyrmions in chiral liquid crystals [35]. That is, to study the flexoelectric polarization effect on skyrmions in liquid crystals. The problem there would be slightly more difficult as the Poisson equation would depend on the divergence of the polarization $\mathbf{P}(\mathbf{n})$ instead of the divergence of the order parameter $\mathbf{n}\in S^{2}$.

### v2

@@ V The Euler-Lagrange equations
$\frac{\textup{d}}{\textup{d}t}\bigg|_{t=0}E_{\textup{exch}}(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\textup{d}^{2}x\,\bm{\varepsilon}\cdot\Delta\mathbf{n}$ (38)
$\frac{\textup{d}}{\textup{d}t}\bigg|_{t=0}E_{\textup{DMI}}(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\textup{d}^{2}x\,\bm{\varepsilon}\cdot(-2\mathbf{d}_{i}\times\partial_{i}\mathbf{n})$ (39)
$\frac{\textup{d}}{\textup{d}t}\bigg|_{t=0}E_{\textup{pot}}(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\textup{d}^{2}x\,\bm{\varepsilon}\cdot(0,0,-h-2Kn_{3}).$ (40)
@@
Denote by $\psi_{t}$ the unique solution of (30) with source ${\rm div}\,\mathbf{n}_{t}$ decaying to 0 at infinity, and $\dot{\psi}=\partial_{t}\psi_{t}|_{t=0}$. It is important to note that, while $\bm{\varepsilon}$ has compact support, neither $\psi=\psi_{0}$ nor $\dot{\psi}$ do: as argued above, they are $1/r$ localized. Care must be taken with boundary terms when computing the variation of $E_{\textup{DDI}}$, therefore. The variation of $E_{\textup{DDI}}$ induced by $\mathbf{n}_{t}$ is
@@
$=\,$ $\frac{1}{\mu}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\psi\Delta\dot{\psi}$
@@
$=\,$ $\frac{1}{\mu}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\psi\Delta\dot{\psi}$
@@
$=\,$ $\frac{1}{\mu}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\psi\Delta\dot{\psi}$ (41)
@@
$=\,$ $\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,{\bm{\varepsilon}}\cdot\bm{\nabla}\psi$
@@
$=\,$ $\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\bm{\varepsilon}\cdot\bm{\nabla}\psi$ (43)
@@
$\frac{\textup{d}}{\textup{d}t}\bigg|_{t=0}E(\mathbf{n}_{t})$ $=$ $\int_{{\mathbb{R}}^{2}}\textup{d}^{2}x\,\bm{\varepsilon}\cdot(\Delta\mathbf{n}-2\mathbf{d}_{i}\times\partial_{i}\mathbf{n}$ (44)
@@ VI Numerical method
We now detail our method for finding magnetization fields $\mathbf{n}$ that simultaneously minimize the energy (32) and solve Poisson’s equation (30). This method is implemented for NVIDIA CUDA architecture and is adapted from a similar non-local method developed to determine (nuclear) skyrmion crystals stabilized by $\omega$-mesons [Leask_Harland_2024, Leask_2024]. Similar methods have also recently been developed to study knot solitons in an extended version of the standard model of particle physics [Hamada_2025] and the back-reaction of the Coulomb force in an extension of the Skyrme model [Gudnason_2025]. We follow the same methodology, separating our algorithm into two stages, detailed as follows.
@@
VI.1 Landau–Lifshitz–Gilbert dynamics
As a consistency check, we verify our numerics by comparing a selcetion of our results with those obtained using Landau–Lifshitz–Gilbert (LLG) dynamics. The dimensionless LLG equation is
$\frac{\textup{d}\mathbf{n}}{\textup{d}\tau}=-\frac{\gamma}{1+\alpha^{2}}\mathbf{n}\times\left(\mathbf{h}_{\textup{eff}}+\alpha\mathbf{n}\times\mathbf{h}_{\textup{eff}}\right)$ (48)
where $\gamma$ is the gyromagnetic ratio, $\alpha$ is the Gilbert damping parameter, the effective magnetic field is
$\mathbf{h}_{\textup{eff}}=\,$ $-\frac{\delta E}{\delta\mathbf{n}}$
$=\,$ $-\Delta\mathbf{n}+2\mathbf{d}_{i}\times\partial_{i}\mathbf{n}+(h+2K\mathbf{n}\cdot\mathbf{e}_{3})\mathbf{e}_{3}+\mathbf{h}_{\textup{demag}},$ (49)
and $\mathbf{h}_{\textup{demag}}=-\bm{\nabla}\psi$ is the demagnetization field. The first term in the LLG equation (48) represents precession and the second term damping. To find static solutions we set $\alpha=1$ and $\gamma=1$. We evolve this using a fourth order Runge–Kutta method. During the evolution we determine the optimal/appropriate magnetostatic potential $\psi$ at each intermediate RK step, i.e. when we compute the slopes $\mathbf{k}_{i}$. We find that the results are consistent, with the LLG method being significantly slower than the arrested Newton flow method (roughly $1\sim 2$ orders of magnitude slower)
@@ VII Results
The micromagnetic parameters employed in this study are chosen to be representative of experimentally realizable chiral ferromagnets that host nanoscale skyrmions. Specifically, we take the exchange stiffness $J=40\,\textup{pJm}^{-1}$, Dzyaloshinskii–Moriya interaction strength $\mathcal{D}=4\,\textup{mJm}^{-2}$, saturation magnetization $M_{s}=580\,\textup{kAm}^{-1}$, and anisotropy constant $K_{m}=0.8\,\textup{MJm}^{-3}$, with no external magnetic field $(B_{\textup{ext}}=0\,\textup{T})$. This gives us the dimensionless coupling constants $K=2$, $\mu=1.057$ and $h=0$. The chosen regime ensures a realistic competition between exchange, Dzyaloshinskii–Moriya, and dipole-dipole energies, allowing all three DMI terms to be compared on equal footing. The inclusion of finite magnetostatic and anisotropy energies further reflects conditions relevant to both thin-film and bulk chiral magnets where dipolar effects significantly influence skyrmion stability.
In particular, these parameters correspond to the energy and length scales typical of soft ferromagnetic multilayers and Heusler-type alloys, such as CoFeB [Tomasello_2014, Wang_2018] and Mn-based compounds [Nayak_2017, Saha_2019], which are widely used as benchmark materials in micromagnetic simulations of chiral spin textures. The length scale $L_{0}=J/\mathcal{D}=10\,\textup{nm}$ defines the intrinsic magnetic length, yielding skyrmion diameters in the experimentally accessible range of $10-100\,\textup{nm}$. The chosen DMI magnitude ensures that chiral twisting competes effectively with exchange and dipolar energies, allowing all three DMI terms – Dresselhaus, Rashba, and Heusler – to produce stable topological textures under comparable conditions.
The saturation magnetization and anisotropy values are consistent with thin-film and bulk ferromagnetic materials where demagnetizing fields are non-negligible, ensuring that the magnetostatic self-interaction plays a physically relevant role [Leonov_2016]. These values thus provide a realistic and well-balanced parameter regime for exploring how dipole-dipole interactions influence skyrmion and antiskyrmion stability in three-dimensional translationally invariant chiral magnets.
For simulations involving isolated magnetic skyrmions, all numerics carried out were simulated on a dimensionless meshgrid of size $10\times 10$ with $1024^{2}$ lattice sites and vacuum boundary conditions. Physically, this corresponds to a $100\textup{nm}\times 100\textup{nm}$ grid with $0.098\textup{nm}\times 0.098\textup{nm}$ lattice spacing. The energy scale for the simulations is $E_{0}=J^{2}/\mathcal{D}=4\times 10^{-19}\,\textup{J}$.
@@ VII.1 Isolated magnetic skyrmions
Dresselhaus Rashba Heusler
Skyrmion Bloch Néel Antiskyrmion
$R_{\textup{Sk}}$ 8.4228.422nm 8.4158.415nm 8.4188.418nm
Table 2: The effect of the DDI on the skyrmion size, $R_{\textup{Sk}}$. The Bloch skyrmion is unchanged, whereas the Néel and anti-skyrmion shrink relative to their size without the DDI ($R_{\textup{Sk}}=8.422$nm).
@@ VII.2 Magnetic skyrmion crystals
Next, we investigate the DDI effect on magnetically ordered crystals. In the absence of DDI, there is no optimal lattice geometry for the parameter set we have chosen – the skyrmions repel one another and want to be infinitely separated. Our method is simpler than that of [Lee_2025]: we do not vary over the lattice geometry. Instead we restrict our geometry to that of two lattice types, an equianharmonic (triangular) lattice and a lemniscatic (square) lattice. In the triangular case, for which the unit cell is a rhombus of side length $L$ and apex angle $\frac{\pi}{3}$, it is convenient to combine unit cells into pairs and work on a rectangular lattice of side lengths $L\times\sqrt{3}L$, with topological charge $|N_{\textup{Sk}}|=2$ per unit cell, see figure 3. For both lattice geometries, we vary the lattice parameter $L$. In all cases, as $L\rightarrow\infty$, the energy per unit topological charge approaches that of the isolated single magnetic skyrmion.
Figure 3: Fields periodic with respect to an equianharmonic lattice of side length $L$ are also periodic with respect to a rectangular lattice of size $L\times\sqrt{3}L$.
@@
In absence of the DDI, there does not exist an optimal lattice size $L_{*}$ which minimizes the energy $E(L)$. The energy is minimized for two infinitely separated skyrmions/antiskyrmions. Consideration of the DDI breaks the energy degeneracy of the crystal configurations, resulting in separate $E$ against $L$ curves for the three DMI terms considered. Qualitatively, the results are similar to those with no dipolar interaction for the Rashba and Dresselhaus DMIs. This is not surprising since isolated skyrmions induce no demagnetizing field in the Dresselhaus case, and no long-range field in the Rashba case: the “charge" distribution $\nabla\cdot\mathbf{n}$ is rotationally symmetric and has total charge 0, so all terms in its multipole expansion vanish, and the demagnetizing field, like $\mathbf{n}$, is exponentially spatially localized.
@@
In more detail, we find that the associated skyrmions maintain their repulsive behaviour and want to be infinitely separated, but the Rashba crystals have higher energy than the Dresselhaus ones. Another key difference is that the energy of the Rashba lattice asymptotically approaches the energy of two well-separated Néel skyrmions at a smaller lattice parameter $L$ than the Dresselhaus lattice does to two Bloch skyrmions. This can be seen in Figures 4 and 5. This is most likely due to the fact that the DDI slightly shrinks the size of the Néel skyrmion, making it more localized than the Bloch skyrmion. Even at low lattice parameter $L$, the DDI has minimal effect on the crystalline structure of Dresselhaus Bloch skyrmions.
However, a rather striking outcome for the Heusler antiskyrmion lattice is observed. First, the antiskyrmions prefer a square lattice arrangement over a triangular one at all lattice sizes $L$. Furthermore, the DDI introduces an attractive force so that there does exist an optimal lattice size, $L_{*}=39$ nm, that minimizes the energy $E(L)$. We have also checked that the square lattice with spacing $L_{*}$ minimizes energy per unit cell with respect to arbitrary variations of the geometry of the period lattice, not just its unit cell size. This calculation is quite sophisticated, and we present the details in an appendix. We are very confident, therefore, that Heusler antiskyrmions bind together to form a square lattice, at the chosen parameter set.
@@
As the lattice size is increased, the energy of the Heusler lattice asymptotically approaches the energy of two infinitely separated antiskyrmions from below, meaning it has positive binding energy $E_{\textup{bind}}=2E_{1}-E_{\textup{latt}}>0$. By contrast, the Rashba and Dresselhaus lattices approach their respective asymptotic energies from above, that is, they have negative binding energies $2E_{1}<E_{\textup{latt}}$. So, it is energetically preferable for the Heusler antiskyrmions to form a crystalline structure, whereas it is preferable for Bloch and Néel skyrmions to be infinitely separated. The results of the effect of the dipolar interaction on the magnetic skyrmion crystals are summarized in Figures 4 and 5.
@@
Figure 5: The effect of the dipolar interaction on square magnetic skyrmion crystals as a function of the lattice size $L$. The unit cell used is of size $L\times L$, which favours a lemniscatic lattice. Initially, two magnetic skyrmions are place in a square arrangement and then are relaxed using the numerical algorithm detailed in the text. It can be seen that including the DDI and its back-reaction lifts the energy degeneracy, resulting in distinct $E(L)$ curves. The Rashba (Néel) and Dresselhaus (Bloch) skyrmion lattices maintain a hexagonal arrangement, whereas the Heusler antiskyrmions relax into one of a square form. An example Heusler antiskyrmion lattice is shown at small $L=20\,\textup{nm}$ and a Rashba Néel skyrmion lattice is shown at a higher $L=30\,\textup{nm}$. Within these subplots is the magnetic charge density $\rho(x,y)$ and the magnetization $\mathbf{n}(x,y)$. Also plotted is the asymptotic energy of two well separated skyrmions for each DMI.
VIII Conclusion
An applied external magnetic field was also considered, but the results were qualitatively the same for our parameter set. However, an external magnetic field in the absence of anisotropies was not considered in this study. For this case, as the applied magnetic field strength is increased, one would expect to see the transition from the spin-stripe domain to a lattice of skyrmions as found in [Kwon_2012].
This method can be adapted to the analogue problem of skyrmions in chiral liquid crystals [Leonov_2014]. That is, to study the flexoelectric polarization effect on skyrmions in liquid crystals. The problem there is slightly more difficult as the Poisson equation depends on the divergence of the polarization $\mathbf{P}(\mathbf{n})$ instead of the divergence of the order parameter $\mathbf{n}\in S^{2}$ [Leask_2025_2].
Appendix A Lattice criticality
@@
We have described a numerical scheme that, for a given fixed period lattice
@@
and topological degree, produces the minimizer among all fields $\mathbf{n}:\mathbb{R}^{2}\rightarrow S^{2}$ periodic with respect to $\Lambda$
@@
$\left.+\frac{1}{2\mu}\psi\Delta\psi\right\}.$ (53)
Here $\Sigma=\mathbb{R}^{2}/\Lambda\equiv T^{2}$ is the torus with period lattice $\Lambda$ (or, in more elementary terms, its unit cell, that is the parallelogram spanned by the vectors $\mathbf{v}_{1},\mathbf{v}_{2}$). We have left the potential terms $V(\mathbf{n})$ arbitrary, since our argument does not rely on the specific potential (4) used in this paper. One should note that such a minimizer is likely to exist for any choice of period lattice $\Lambda$. If this solution is to have any physical significance, it should be a local minimizer of $E_{\textup{cell}}$ with respect to variations of the period lattice $\Lambda$ also. Previous studies of soliton crystals in other systems show that, in general, the energetically optimal period lattice may have unexpectedly low symmetry, being neither square nor triangular [Speight_Winyard, Leask_2022], so this issue is not negligible. In this appendix, we formulate this criticality condition as a set of explicit integral constraints, suitable for numerical verification, adapting ideas introduced in [Speight_2014].
@@
$(\partial_{1}^{2}+\partial_{2}^{2})\widetilde{\psi}=-\mu(\partial_{1}\widetilde{\mathbf{n}}+\partial_{2}\widetilde{\mathbf{n}}_{2}).$ (54)
@@
[…]
@@
$\frac{\textup{d}}{\textup{d}t}\bigg|_{t=0}E_{\textup{pot}}^{\textup{cell}}(A(t))$ $=$ $0.$ (70)